OpenStudy (anonymous):

Find intervals on which each function is continuous...again sorry :( i'll learn, i swear xD

OpenStudy (anonymous):

\[f(x)=\frac{ x ^{2} }{ 2x+4 }\]

OpenStudy (anonymous):

my answer was all real #?

OpenStudy (anonymous):

graph it and look for places where there is a large jump in y between points or where points are missing

OpenStudy (anonymous):

I would look at where the denominator is zero and see if that point exists. if it doesn't, the function cannot be continuous there.

OpenStudy (anonymous):

ok so -2 would make the denominator be 0.

OpenStudy (anonymous):

does that mean there is a hole there?

OpenStudy (anonymous):

Terenz help!!! please

terenzreignz (terenzreignz):

Rational functions like this one are continuous everywhere in their domain, so your best bet is finding a point that is not in its domain.

OpenStudy (anonymous):

Sorry, i don't really get it. i'm trying to graph it, what kind of graph is it?

terenzreignz (terenzreignz):

There is no need to graph. Just look for the points (or x-values) for which this function has no existent value. Points of discontinuity of rational functions are points for which the denominator becomes zero.

OpenStudy (anonymous):

So it's -2?

terenzreignz (terenzreignz):

Yes. That seems to be the only point.

terenzreignz (terenzreignz):

But you are not asked for points of discontinuity, rather, you are asked for intervals for which this function is continuous... got that handled?

OpenStudy (anonymous):

and the intervals would be..(-infinity,2)(2,infinity)

terenzreignz (terenzreignz):

<ehem>

OpenStudy (anonymous):

-infinty<-2>infinity?

terenzreignz (terenzreignz):

\[\Large (-\infty,-2)\cup(-2,\infty)\]

OpenStudy (anonymous):

oh..sorry yea

OpenStudy (anonymous):

Thank you

terenzreignz (terenzreignz):

No problem.